# Density Frankl–Rödl on the Sphere

### v1

@@ 3.2 Reverse Hypercontractivity
$\mathop{\mathbb{E}\/}_{x\cdot y=e^{-t}}[f(x)g(y)]=\langle f,A_{t}g\rangle\geq\|f\|_{p}\|g\|_{p}-O_{t}(n^{-2/3})\|f\|_{2}\|g\|_{2},$
@@
$\mathop{\mathbb{E}\/}_{x\cdot y=e^{-t}}[f(x)g(x)]=\langle f,A_{t}g\rangle\geq\langle f,P_{t}g\rangle-O_{t}(n^{-2/3})\|f\|_{2}\|g\|_{2}.$
Let $p^{\prime}=\frac{p}{p-1}$ be the Hölder conjugate of $p$, so that $11/p+1/p^{\prime}=$ 1. Applying the reverse Hölder inequality (see [HLP52], Theorem 13), we obtain
@@ 4.1 Inductive Configurations
Let $(v_{1},\cdots,v_{k})$ be a configuration on ${\mathbb{S}^{n-1}}$, where $v_{i}\in{\mathbb{S}^{n-1}}$ are distinct vectors satisfying $\left\langle{v_{i},v_{j}}\right\rangle=r_{i}$ for $i<j$. Specifically, we consider configurations whose covariance matrix has the form
@@
Start with any vector $v_{1}$ and a choice of $-1<r_{1}<1$. Choose $v_{2}$ on the $(n-2)$-dimensional subsphere $S_{v_{1},r_{1}}=\{x\in{\mathbb{S}^{n-1}}:\left\langle{x,v_{1}}\right\rangle=r_{1}\}$. Next, select $r_{2}$ such that the intersection $S_{v_{1},r_{1}}\cap S_{v_{2},r_{2}}$ is nonempty, and choose $v_{3}$ on this $(n-3)$-dimensional subsphere. Proceed inductively: for each $v_{i}$, choose it on the $(n-i)$-dimensional subsphere $\bigcap_{j<i}S_{v_{j},r_{j}}$.
Throughout the paper, we exclude a special case of inductive configuration–configurations where $v_{k}-v_{k-1}$ is the diameter of $\bigcap_{j<k-1}S_{v_{j},r_{j}}$–for reasons that will become clear in the later proof. Notably, both the length of $v_{k}-v_{k-1}$ and the diameter of $\bigcap_{j<k-1}S_{v_{j},r_{j}}$ are independent of the ambient dimension $n$, so this is a constraint on the configuration $R$ itself.
@@ 4.2 Configuration with Pairwise Orthogonal Vectors
Let $\Delta_{k}(n,0):=\Delta(n,I_{k})=\{(x_{1},\cdots,x_{k})\in({\mathbb{S}^{n-1}})^{k}:\left\langle{v_{i},v_{j}}\right\rangle=0,\ \forall i\neq j\}$ denote the set of all $k$-tuples of pairwise orthogonal vectors in ${\mathbb{S}^{n-1}}$.
@@
$\mathop{\bf Pr\/}_{(x_{1},\cdots,x_{k})\in\Delta_{k}(n,0)}(x_{1}\in A,\cdots,x_{k}\in A)\geq\Omega\left(\sigma(A)^{k}-\exp(-cn^{1/3})\sigma(A)^{k-1}\right),$
where $c>0$ is a universal constant.
@@
[…]
When $k>2$, assume the claim holds for $k-1$. Then:
@@
[…]
@@
[…]
@@
[…]
@@ 4.3 Main Results
Theorem 4.3.
@@
$\mathop{\bf Pr\/}_{x\cdot y=r}(x\in A,\ y\in B)\geq(\sigma(A)\sigma(B))^{1/(1-r^{4})}-O(n^{-2/3})\sqrt{\sigma(A)\sigma(B)}.$
@@
$\geq\|f\|_{p}\|g\|_{p}-O(n^{-2/3})\|f\|_{2}\|g\|_{2}$
$=(\sigma(A)\sigma(B))^{1/(1-r^{4})}-O(n^{-2/3})\sqrt{\sigma(A)\sigma(B)}\ .\qed$
@@
Using Theorem 4.3, we can show that many such good vectors must exist:
Proposition 4.4.
@@
Then
$\sigma(A_{\mathrm{good}})\geq\frac{1}{2}\sigma(A)^{2/(1-r^{4})}-O_{r}(n^{-2/3})\sigma(A).$
@@
$\mathop{\mathbb{E}\/}_{x}[\sigma_{x,r}(A\cap S_{x,r})\mid x\in A]=\mathop{\bf Pr\/}_{x\cdot y=r}(y\in A\mid x\in A)\geq\sigma(A)^{(1+r^{4})/(1-r^{4})}-O_{r}(n^{-2/3}).$
@@
$\sigma(A_{\mathrm{good}})=\mathop{\bf Pr\/}(x\in A_{\mathrm{good}}\mid x\in A)\cdot\sigma(A)\geq\frac{1}{2}\sigma(A)^{2/(1-r^{4})}-O_{r}(n^{-2/3})\sigma(A).$
@@
Then for any measurable set $A\subseteq\mathbb{S}^{n-1}$ with $\sigma(A)\geq\omega(n^{-2/3})$, the probability that a uniformly random tuple $(x_{1},\dots,x_{k})\in\Delta(n,R)$ lies entirely in $A$ satisfies
@@
and $f_{c}(r)=\frac{r-c^{2}}{1-c^{2}}$.
@@
We proceed by induction on $k$. The base case $k=2$ follows directly from Theorem 4.3. The condition $v_{2}-v_{1}\neq\operatorname{diam}({\mathbb{S}^{n-1}})$ ensures $r=\left\langle{v_{1},v_{2}}\right\rangle>-1$.
Assume the result holds for all configurations of size $<k$, and let $c_{1}=r_{1}$. Define
@@
Given $x_{1}$, the event $x_{2}\in A,\cdots,x_{k}\in A$ is equal to $y_{2}\in A\cap S_{x_{1},c_{1}},\cdots,y_{k}\in A\cap S_{x_{1},c_{1}}$. Conditioned on $x_{1}\in A_{\mathrm{good}}$, we have $\sigma_{x_{1},c_{1}}(A\cap S_{x_{1},c_{1}})\geq\frac{1}{2}\sigma(A)^{(1+c_{1}^{4})/(1-c_{1}^{4})}$, so by the inductive hypothesis:
@@
Multiplying the two bounds gives
@@
$v_{k}-v_{k-1}\neq\operatorname{diam}\left(\bigcap_{j<k-1}S_{v_{j},r_{j}}\right)$
is equivalent to $c_{k-1}\neq-1$, which is necessary for applying Theorem 4.3. For a given configuration $(v_{1},\dots,v_{k})$, the quantity $c_{i}$ corresponds to the inner product between $v_{i}$ and later vectors $v_{j}$, after projecting to the $(n-i)$-subsphere $\bigcap_{l<i}S_{v_{l},r_{l}}$ and normalizing. Since $v_{1},\dots,v_{k}$ are distinct, we have $-1<c_{i}<1$ for all $i\leq k-2$ and $c_{k-1}<1$. The only forbidden case is $c_{k-1}=-1$, which occurs precisely when $v_{k}-v_{k-1}$ equals the diameter of the intersection $\bigcap_{j<k-1}S_{v_{j},r_{j}}$.
@@
Fix any $r\in\left(-\frac{1}{k-1},1\right)$. For any measurable set $A\subseteq\mathbb{S}^{n-1}$ with $\sigma(A)\geq\omega(n^{-2/3})$, the probability that a uniformly random tuple $(x_{1},\dots,x_{k})\in\Delta_{k}(n,r)$ lies entirely in $A$ satisfies
$\mathop{\bf Pr\/}_{(x_{1},\dots,x_{k})\in\Delta_{k}(n,r)}(x_{1}\in A,\dots,x_{k}\in A)\geq\Omega_{k,r}(\sigma(A)^{C_{R}}),$
as $n\to\infty$. The constant $C_{R}$ is given by
@@ 5 Open Problems
Our density version of the Frankl–Rödl theorem on the sphere (Theorem 4.3) requires the set $A\subseteq\mathbb{S}^{n-1}$ to have spherical measure $\sigma(A)>\omega(n^{-2/3})$ in order to guarantee a non-trivial lower bound. In contrast, the classical Frankl–Rödl theorem asserts that for any $-1<r<1$, there exists a constant $\epsilon=\epsilon(r)>0$ such that
@@
for all $-1<r<1$, even when $\sigma(A)$ is exponentially small in $n$.
Another direction concerns the class of configurations we consider. Our density result applies to inductive configurations, but Matoušek and Rödl [MR95] showed that for any configuration $P$ with circumradius less than 1, the probability
$\mathop{\bf Pr\/}_{(x_{1},\ldots,x_{k})\in P}(x_{i}\in A\ \forall i)>0$
is positive for any set $A\subseteq\mathbb{S}^{n-1}$ with constant density $\sigma(A)>\epsilon(P)$. It remains open whether one can establish a density version of this theorem—that is, to derive quantitative lower bounds on the probability in terms of $\sigma(A)$.

### v3

@@ 3.2 Reverse Hypercontractivity
$\mathop{\mathbb{E}\/}_{x\cdot y=e^{-t}}[f(x)g(y)]=\langle f,A_{t}g\rangle\geq\|f\|_{p}\|g\|_{p}-O_{t}\left(\frac{1}{n}\right)\|f\|_{2}\|g\|_{2},$
@@
$\mathop{\mathbb{E}\/}_{x\cdot y=e^{-t}}[f(x)g(x)]=\langle f,A_{t}g\rangle\geq\langle f,P_{t}g\rangle-O_{t}\left(\frac{1}{n}\right)\|f\|_{2}\|g\|_{2}.$
Let $p^{\prime}=\frac{p}{p-1}$ be the Hölder conjugate of $p$, so that $1/p+1/p^{\prime}=1$. Applying the reverse Hölder inequality (see [HLP52], Theorem 13), we obtain
@@ 4.1 Inductive Configurations
Let $(v_{1},\cdots,v_{k})$ be a configuration on $\mathbb{S}^{n-1}$}, where $v_{i}\in{\mathbb{S}^{n-1}}$ are distinct vectors satisfying $\left\langle{v_{i},v_{j}}\right\rangle=r_{i}$ for $i<j$. Specifically, writing $v_{i}$ as column vectors, we consider configurations whose covariance matrix has the form
@@
Start with any vector $v_{1}$ and a choice of $1<r_{1}<1$. Choose $v_{2}$ on the (n2)(n-2)-subsphere $S_{v_{1},r_{1}}=\{x\in{\mathbb{S}^{n-1}}:\left\langle{x,v_{1}}\right\rangle=r_{1}\}$. Next, select $r_{2}$ such that the intersection $S_{v_{1},r_{1}}\cap S_{v_{2},r_{2}}$ is nonempty, and choose $v_{3}$ on this (n3)(n-3)-subsphere. Proceed inductively: for each $v_{i}$, choose it on the (ni)(n-i)-subsphere $\bigcap_{j<i}S_{v_{j},r_{j}}$.
Throughout the paper, we exclude a special case of inductive configuration–configurations where the length of $v_{k}-v_{k-1}$ is the diameter of $\bigcap_{j<k-1}S_{v_{j},r_{j}}$—for reasons that will become clear in the later proof. Notably, both the length of $v_{k}-v_{k-1}$ and the diameter of $\bigcap_{j<k-1}S_{v_{j},r_{j}}$ are independent of the ambient dimension $n$, so this is a constraint on the configuration $R$ itself.
@@ 4.2 Configuration with Pairwise Orthogonal Vectors
Let $\Delta_{k}(n,0):=\Delta(n,I_{k})=\{(x_{1},\cdots,x_{k})\in({\mathbb{S}^{n-1}})^{k}:\left\langle{x_{i},x_{j}}\right\rangle=0,\ \forall i\neq j\}$ denote the set of all $k$-tuples of pairwise orthogonal vectors in $\mathbb{S}^{n-1}$}. Recall that $S_{x,r}$ denotes the (n2)(n-2)-subsphere $\{y\in\mathbb{S}^{n-1}:x\cdot y=r\}$ and $\sigma_{x,r}$ is the uniform probability measure on $S_{x,r}$.
@@
$\mathop{\bf Pr\/}_{(x_{1},\cdots,x_{k})\in\Delta_{k}(n,0)}(x_{1}\in A,\cdots,x_{k}\in A)\geq\Omega_{k}\left(\sigma(A)^{k}-C_{k}\exp(-cn^{1/3})\sigma(A)^{k-1}\right),$
where $c>0$ is a universal constant and $C_{k}$ is a constant depending only on $k$.
@@
$\geq{}$ $\mathop{\bf Pr\/}_{(x_{1},x_{2})\in\Delta_{2}(n,0)}(x_{1}\in A\cap B,x_{2}\in A)$
$={}$ $\mathop{\bf Pr\/}(x_{1}\in A\cap B)\cdot\mathop{\bf Pr\/}(x_{2}\in A\mid x_{1}\in A\cap B)$
$={}$ $\mathop{\bf Pr\/}(x_{1}\in A\cap B)\cdot\mathop{\mathbb{E}\/}[\sigma_{x_{1},0}(A\cap S_{x_{1},0})\mid x_{1}\in A\cap B]$
$\geq{}$ $\left(\sigma(A)-C\exp(-cn^{1/3})\right)\cdot 0.9\sigma(A)$
$\geq{}$ $\Omega\left(\sigma(A)^{2}-C\exp(-cn^{1/3})\sigma(A)\right)$
When $k>2$, assume the claim holds for k1k-1. Then:
@@
$\geq{}$ $\mathop{\bf Pr\/}(x_{1}\in A\cap B)\cdot\mathop{\bf Pr\/}(x_{2},\cdots,x_{k}\in A\mid x_{1}\in A\cap B)$
$={}$ $\mathop{\bf Pr\/}(x_{1}\in A\cap B)\cdot\mathop{\bf Pr\/}(x_{2},\cdots,x_{k}\in A\cap S_{x_{1},0}\mid x_{1}\in A\cap B).$
@@
$\geq{}$ $\Omega_{k}\left((0.9\sigma(A))^{k-1}-C_{k}\exp(-cn^{1/3})(0.9\sigma(A))^{k-2}\right)$
$\geq{}$ $\Omega_{k}\left(\sigma(A)^{k-1}-C_{k}\exp(-cn^{1/3})\sigma(A)^{k-2}\right).$
@@
$\geq{}$ $\Omega_{k}\left(\sigma(A)^{k-1}-C_{k}\exp(-cn^{1/3})\sigma(A)^{k-2}\right)\cdot\left(\sigma(A)-C\exp(-cn^{1/3})\right)$
$\geq{}$ $\Omega_{k}\left(\sigma(A)^{k}-C^{\prime}_{k}\exp(-cn^{1/3})\sigma(A)^{k-1}\right),$
@@ 4.3 Main Results
We first prove a two-set version of the density sphere Frankl-Rödl theorem which will be later used for induction on inductive configurations.
Theorem 4.3 (Density Sphere Frankl-Rödl).
@@
$\mathop{\bf Pr\/}_{x\cdot y=r}(x\in A,\ y\in B)\geq(\sigma(A)\sigma(B))^{1/(1-r^{4})}-O_{r}\left(\frac{1}{n}\right)\sqrt{\sigma(A)\sigma(B)}.$
@@
$\geq\|f\|_{p}\|g\|_{p}-O_{r}\left(\frac{1}{n}\right)\|f\|_{2}\|g\|_{2}$
$=(\sigma(A)\sigma(B))^{1/(1-r^{4})}-O_{r}\left(\frac{1}{n}\right)\sqrt{\sigma(A)\sigma(B)}\ .\qed$
@@
Using Theorem 4.3, we can show that many such good vectors must exist.
Proposition 4.4.
The set of good vectors satisfies
$\sigma(A_{\mathrm{good}})\geq\frac{1}{2}\sigma(A)^{2/(1-r^{4})}-O_{r}\left(\frac{1}{n}\right)\sigma(A).$
@@
$\mathop{\mathbb{E}\/}_{x}[\sigma_{x,r}(A\cap S_{x,r})\mid x\in A]=\mathop{\bf Pr\/}_{x\cdot y=r}(y\in A\mid x\in A)\geq\sigma(A)^{(1+r^{4})/(1-r^{4})}-O_{r}\left(\frac{1}{n}\right).$
@@
$\sigma(A_{\mathrm{good}})=\mathop{\bf Pr\/}(x\in A_{\mathrm{good}}\mid x\in A)\cdot\sigma(A)\geq\frac{1}{2}\sigma(A)^{2/(1-r^{4})}-O_{r}\left(\frac{1}{n}\right)\sigma(A)\ .\qed$
@@
$\left\lVert v_{k}-v_{k-1}\right\rVert_{2}\neq\operatorname{diam}\left(\bigcap_{j<k-1}S_{v_{j},r_{j}}\right).$
Then for any measurable set $A\subseteq\mathbb{S}^{n-1}$ with $\sigma(A)\geq\omega_{R}(n^{-\epsilon_{R}})$, the probability that a uniformly random tuple $(x_{1},\dots,x_{k})\in\Delta(n,R)$ lies entirely in $A$ satisfies
@@
and $f_{c}(r)=\frac{r-c^{2}}{1-c^{2}}$. The constant $\epsilon_{R}$ is given by
$\epsilon_{R}=\prod_{i=1}^{k-1}\frac{1-c_{i}^{4}}{1+c_{i}^{4}}.$
@@
We proceed by induction on $k$. The base case $k=2$ follows directly from Theorem 4.3. The condition $v_{2}-v_{1}\neq\operatorname{diam}({\mathbb{S}^{n-1}})$ ensures $r=\left\langle{v_{1},v_{2}}\right\rangle>-1$ and the condition $\sigma(A)\geq\omega_{R}(n^{-\epsilon_{R}})$ ensures that the error term goes to 0.
Assume the result holds for all configurations of size <k<k, and let $c_{1}=r_{1}$. Define
@@
Given $x_{1}$, the event $x_{2}\in A,\cdots,x_{k}\in A$ is equal to $y_{2}\in A\cap S_{x_{1},c_{1}},\cdots,y_{k}\in A\cap S_{x_{1},c_{1}}$. Conditioned on $x_{1}\in A_{\mathrm{good}}$, we have $\sigma_{x_{1},c_{1}}(A\cap S_{x_{1},c_{1}})\geq\frac{1}{2}\sigma(A)^{(1+c_{1}^{4})/(1-c_{1}^{4})}\geq\omega_{R^{\prime}}(n^{-\epsilon_{R^{\prime}}})$, so by the inductive hypothesis:
@@
Multiplying the two bounds give
@@
$\left\lVert v_{k}-v_{k-1}\right\rVert_{2}\neq\operatorname{diam}\left(\bigcap_{j<k-1}S_{v_{j},r_{j}}\right)$
is equivalent to $c_{k-1}\neq-1$, which is necessary for applying Theorem 4.3. For a given configuration $(v_{1},\dots,v_{k})$, the quantity $c_{i}$ corresponds to the inner product between $v_{i}$ and later vectors $v_{j}$, after projecting to the (ni)(n-i)-subsphere $\bigcap_{l<i}S_{v_{l},r_{l}}$ and normalizing. Since $v_{1},\dots,v_{k}$ are distinct, we have $1<c_{i}<1$ for all $i\leq k-2$ and $c_{k-1}<1$. The only forbidden case is $c_{k-1}=-1$, which occurs precisely when $v_{k}-v_{k-1}$ equals the diameter of the intersection $\bigcap_{j<k-1}S_{v_{j},r_{j}}$.
@@
Fix any $r\in\left(-\frac{1}{k-1},1\right)$. For any measurable set $A\subseteq\mathbb{S}^{n-1}$ with $\sigma(A)\geq\omega_{k,r}(n^{-\epsilon_{k,r}})$, the probability that a uniformly random tuple $(x_{1},\dots,x_{k})\in\Delta_{k}(n,r)$ lies entirely in $A$ satisfies
$\mathop{\bf Pr\/}_{(x_{1},\dots,x_{k})\in\Delta_{k}(n,r)}(x_{1}\in A,\dots,x_{k}\in A)\geq\Omega_{k,r}(\sigma(A)^{C_{k,r}}),$
as $n\to\infty$. The constants $C_{k,r},\ \epsilon_{k,r}$ are given by
$C_{k,r}=\sum_{i=1}^{k-1}\frac{2}{1-c_{i}^{4}}\prod_{j=1}^{i-1}\frac{1+c_{j}^{4}}{1-c_{j}^{4}}\qquad\epsilon_{k,r}=\prod_{i=1}^{k-1}\frac{1-c_{i}^{4}}{1+c_{i}^{4}},$
@@ 5 Open Problems
Our density version of the Frankl–Rödl theorem on the sphere (Theorem 4.3) requires the set $A\subseteq\mathbb{S}^{n-1}$ to have spherical measure $\sigma(A)>\omega\left(\frac{1}{n}\right)$ in order to guarantee a non-trivial lower bound. In contrast, the classical Frankl–Rödl theorem asserts that for any 1<r<1-1<r<1, there exists a constant $\epsilon=\epsilon(r)>0$ such that
@@
for all 1<r<1-1<r<1, even when $\sigma(A)$ is exponentially small in $n$.
Another direction concerns the class of configurations we consider. Our density result applies to inductive configurations, but Matoušek and Rödl [MR95] showed that fix constant $c$ and for any configuration $P$ with circumradius less than 1, any $c$ coloring of $\mathbb{S}^{n-1}$} contains a monochromatic copy of $P$. It remains open whether one can establish a density version of this theorem—that is, to show
$\mathop{\bf Pr\/}_{(x_{1},\ldots,x_{k})\in P}(x_{i}\in A\ \forall i)>\epsilon(\sigma(A))$
for any set $A\subseteq\mathbb{S}^{n-1}$ with constant density $\sigma(A)>\epsilon(P)$ where $\epsilon(\sigma(A))$ is a lower bound on the probability in terms of $\sigma(A)$.